what is the slope of lines passing through (1,-1) and (3,5)
Answer by venugopalramana(3286)

what is the slope of lines passing through (1,-1)=x1,y1 and (3,5)=x2,y2
slope = (y2-y1)/(x2-x1)
= (5+1)/(3-1)=3